# Constraints on the anomalous Higgs boson couplings in $Z\gamma\gamma$ production at muon collider

### v1

@@ Abstract
We investigate the sensitivities on the Wilson coefficients of dimension-six operators associated with the anomalous $H\gamma\gamma$, $H\gamma Z$ and $HZZ$ vertices through the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ at future muon collider. Signal events involving Higgs-gauge boson interactions and relevant backgrounds events at the muon collider designed with center-of-mass energy of 10 TeV and integrated luminosity of 10 ab-1 are generated in Madgraph within the framework of the model-independent Standard Model effective field theory and detector effects are also included with corresponding detector cards in Delphes. The limits at 95% C.L. on the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are reported to be [-0.0052; 0.0040] and [-0.0034; 0.0023], respectively, and compared with the experimental and phenomenological limits.
@@ I Introduction
With the discovery of the Higgs particle Aad:2012les ; Chatrchyan:2012les , intensive studies have been done to reveal the properties of the Higgs particle at colliders, both through the Standard Model (SM) predictions and scenarios beyond the SM. Despite the discovery of the Higgs boson, the Higgs sector’s structure is still unclear. Despite many successful predictions of the SM, its clear discrepancies with some experimental observations and its inability to explain some major problems such as dark matter, dark energy, hierarchy problem, strong CP problem, etc. encourage us to search for indirect signatures of new physics.
New physics extensions beyond the SM are necessary to address unresolved open questions like the observed baryon-antibaryon asymmetry of the Universe, even if studies Aad:2016saf ; Aad:2020czq ; Aad:2020jkn ; Sirunyan:2020wxc ; Tumasyan:2022ssx indicate that the currently discovered Higgs boson is a CP-even scalar with CP-conserving interactions and is in accord with the SM predictions. In cosmology and particle physics, one of the biggest puzzles is the Baryon Asymmetry of the Universe (BAU). If three Sakharov conditions Sakharov:1967enp are met, the mechanism of electroweak baryogenesis (EWBG) Kuzmin:1985yla can explain the matter-antimatter asymmetry. These conditions are: (i) baryon number violation, (ii) C and CP violation, and (iii) deviation from the thermodynamic equilibrium.
The quantity of CP violation in the SM caused by the Cabibbo-Kobayashi-Maskawa (CKM) matrix Cabibbo:1963les ; Kobayashi:1973hqw is insufficient to quantitatively reproduce the BAU Gavela:1994ogb ; Morrissey:2012ydf if we concentrate on the second of these three conditions. Extended Higgs sectors that offer more sources of CP violation can solve this issue. In particle physics, one crucial area of study for future experiments is the search for CP violation. One of the prerequisites for baryogengesis is CP violation, and quark flavor physics which is in accord with the CKM mechanism in the SM Kobayashi:1973hqw provides the only experimental proof of CP violation to date. Therefore, an intriguing way to search for a new mechanism is to look for CP violation in the interactions between the Higgs boson and gauge bosons. A major goal of experimental measurements is to restrict the certain range of CP-violating parameters or to find nonzero CP violation in the Higgs boson interactions. Future high-energy physics experiments have special qualifications to test for CP violation in the Higgs boson interactions. Since CP violation impacts are negligible in the SM, they offer great testing opportunities to compare the performance of future facilities. The studies of Higgs boson phenomenology have come into prominence in searching for the SM and new physics effects at current and future colliders. The search for precise measurements of Higgs boson couplings at multi-TeV muon colliders has recently become more important in the study of physics beyond the SM Chiesa:2020jer ; Han:2021opj ; Buttazzo:2021uha ; Han:2021egh ; Forslund:2022hjw ; Celada:2024tgv ; Chiesa:2024asx ; Cassidy:2024uhn ; Dermisek:2024qaz ; Spor:2024ghq ; Forslund:2024yhz ; Gurkanli:2024wcv ; Chen:2022ygc ; Li:2024pmz ; Spor:2024yln . The presence of theories beyond the SM that can predict effects of CP violation will have significant outcomes for particle physics in the future.
@@ II Effective operators and Higgs-gauge boson interactions
A new physics model with additional fields and interactions is needed to make progress in particle physics. In order to understand the new physics in light of experimental data, it is necessary to establish the relationship between the observables at the electroweak scale and physics beyond the SM at a relatively high scale. In order to describe new physics at high mass scales, effective field theory (EFT) presents a model-independent way. EFTs are a tool to bridge the gap between scales and reveal the indirect the new physics effects from experimental data.
The EFT is only valid up to a particular energy scale ($\Lambda$); in order for the theory to be practical, it must be above the energy scale ($E$) that is immediately accessible experimentally. In other words, when $\Lambda\gg E$, the EFT gives a good approximation. In the EFT framework, the effects of unobserved states, which are predicted to arise at energies bigger than an effective scale determined by the $W$-boson mass $m_{W}$, are parameterized using high-dimensional operators for interactions beyond the SM.
One of the most effective ways to measure the effects of the new physics is to use higher-dimensional effective operators consisting of SM fields that obey the SM gauge symmetry. These are named the Standard Model Effective Field Theory (SMEFT) operators. In this approximation, additional dimension-six operators are added to the SM Lagrangian to ensure conservation of lepton and baryon numbers, and dimension-eight and higher operators are neglected for further suppression. The SMEFT operators can be expressed in a theoretical context in the Strongly Interacting Light Higgs (SILH) basis.
An effective theory of a light composite Higgs boson, which is responsible for the electroweak symmetry breaking and arises as the pseudo-Goldstone boson from a strongly interacting sector, reveals the SILH basis Tosciri:2021yhb . The effective Lagrangian of the $H\gamma\gamma$, $HZZ$ and $H\gamma Z$ couplings in the SILH basis Giudice:2007ops is described as the sum of the Lagrangian involving the CP-conserving and CP-violating dimension-six operators and the dimension-four SM Lagrangian as follows:
@@
where ${\overline{c}_{i}}$ are the Wilson coefficients normalized with the new physics scale $\Lambda$ determined by the $W$-boson mass $m_{W}$ and ${\cal O}_{i}$ are the dimension-six operators. The effective Lagrangian with the CP-conserving and CP-violating operators corresponding to the second and third terms in Eq. (1), respectively, are shown below Alloul:2014hws
@@
where $h$, $F_{\mu\nu}$ and $Z_{\mu\nu}$ are the Higgs-boson field, the field strength tensors of photon and $Z$-boson, respectively. $\widetilde{F}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ and $\widetilde{Z}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}Z_{\rho\sigma}$ are the dual field strength tensors. The relationships among the Lagrangian parameters in the mass basis in Eq. (4) and the gauge basis in Eqs. (2-3) are shown below:
@@
We have performed analyses of the effects of dimension-six operators on Higgs-gauge boson couplings in MadGraph5__\__aMC@NLO Alwall:2014cvc based on Monte Carlo simulations using FeynRules Alloul:2014tfc . In the Higgs effective Lagrangian model file of the FeynRules, the SILH Lagrangian is included along with the Wilson coefficients for Higgs boson interactions.
We investigate the potential of anomalous Higgs-gauge boson couplings on the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$ at the muon collider and the sensitivity study of coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$ and $\widetilde{c}_{HW}$ in anomalous $H\gamma\gamma$, $HZZ$ and $H\gamma Z$ vertices. The stage of the muon collider with a center-of-mass energy of 10 TeV and an integrated luminosity of 10 ab-1 is considered Accettura:2023uhz . By combining high-energy collisions in a clean experimental environment, the multi-TeV muon collider creates unique potential for Higgs studies by providing both precise measurements and high-energy reach in a single machine. Therefore, a $10+$ TeV lepton collider that goes beyond the precision-energy dilemma is the ideal machine. Since electron-based colliders using current technology are unable to reach this energy scale, we are turning to muon colliders to probe this scale with leptons Black:2024edf .
The Feynman diagrams of the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ are shown in Fig. 1. The three diagrams in the top row are pure SM background processes, while the other six diagrams present signal processes, including new physics contributions from the anomalous $H\gamma\gamma$ (black dot), $H\gamma Z$ (red dot) and $HZZ$ (blue dot) vertices.
@@
The total cross-sections of the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ as a function of coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$ and $\widetilde{c}_{HW}$ at the muon collider are given in Fig. 2. The transverse momentum $p_{T}^{\gamma}>10$ GeV and the pseudo-rapidity $|\eta^{\gamma}|<2.5$ for the photons on the final state particles are applied to determine the total cross-sections in Fig. 2. While the other coefficients are fixed to zero, the coefficient under consideration is variable each time at the calculation method of total cross-sections. The SM cross-section including the contributions from the top three Feynman diagrams in Fig. 1 corresponds to the point $\overline{c}_{\gamma}=\overline{c}_{HB}=\overline{c}_{HW}=\widetilde{c}_{\gamma}=\widetilde{c}_{HB}=\widetilde{c}_{HW}=0$ in Fig. 2.
@@ III Analysis of Signal and Background Events
We describe the cut-based analysis and the detector simulation to determine the limits on anomalous Higgs-gauge boson couplings for the $H\gamma\gamma$, $HZZ$ and $H\gamma Z$ vertices at the muon collider. The process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$ with nonzero coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{HW}$ is considered as signal process that comprises the SM contribution as well as interference between effective couplings and SM contributions. Two relevant backgrounds are considered; the SM contribution from the same final state of the signal process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$, where $\ell$ is a charged lepton such as electron $(e^{\pm})$ and muon $(\mu^{\pm})$, is considered as the main background process, and the process $\mu^{-}\mu^{+}\rightarrow\tau\tau\gamma\gamma$ is considered as the minor background process.
@@
In MadGraph5__\__aMC@NLO, 500k events are generated for each coefficient value of both the signal and the background processes. These events, which involve initial and final parton shower, fragmentation and decay, are processed through the Pythia 8.3 Bierlich:2022uzx . The Delphes 3.5.0 package Favereau:2014wfb with the muon collider configuration card is used to simulate the detector responses and all events are analyzed with ROOT 6 Brun:1997gqa .
The $\ell\ell\gamma\gamma$ channel is subjected to independent applications of the preselection and kinematic cuts in both signal and background processes to differentiate the signal from the relevant backgrounds. When identifying kinematic cuts, it is essential to consider that the signal has a clear signature with respect to the backgrounds. For the signal and the relevant backgrounds, we therefore take into account different kinematic distributions based on the cuts. Preselection is defined in the study for events with two isolated photons in the $Z\gamma\gamma$ production and an opposite-sign same flavor dilepton in the decay of the $Z$-boson and this preselection is labeled with Cut-0.
@@
According to their transverse momentum, the leading and sub-leading charged leptons ($\ell_{1}$ and $\ell_{2}$) and photons ($\gamma_{1}$ and $\gamma_{2}$) are arranged as follows: $p_{T}^{\ell_{1}}>p_{T}^{\ell_{2}}$ and $p_{T}^{\gamma_{1}}>p_{T}^{\gamma_{2}}$, respectively. Fig. 3 shows the distributions of the transverse momentum of the leading charged lepton and the leading photon for the signal and the relevant background processes at the muon collider. It can be seen that the signal can be distinguished from the backgrounds by $p_{T}^{\ell_{1}}>50$ GeV and $p_{T}^{\gamma_{1}}>50$ GeV. The transverse momentum and pseudo-rapidity distributions of the leading and sub-leading charged lepton for the events are determined as $p_{T}^{\ell_{1}}>50$ GeV, $p_{T}^{\ell_{2}}>10$ GeV and $|\eta^{\ell_{1,2}}|\leq 2.5$, respectively, and are labeled Cut-1. However, the distributions of transverse momentum and pseudo-rapidity of the leading and sub-leading photons for the events are determined as $p_{T}^{\gamma_{1}}>50$ GeV, $p_{T}^{\gamma_{2}}>10$ GeV and $|\eta^{\gamma_{1,2}}|\leq 2.5$, respectively, and are labeled Cut-2.
@@
Figure 3: The distributions of transverse momentum of the leading charged lepton (left) and the leading photon (right) for signal and relevant background processes at the muon collider.
@@
A clear difference between signal and background is observed in the distribution of the azimuthal angle between the leading and sub-leading photons and charged leptons in Fig. 4 and in the distribution of the minimum distance between the leading and sub-leading photons and charged leptons in Fig. 5. For photons and charged leptons to be well separated in phase space leading to their identification as distinct objects in detector, kinematic cuts are needed such as $\Delta\phi_{\gamma\gamma}<0.8$ and $\Delta\phi_{\ell\ell}<1.8$ at Cut-3 and $\Delta R_{\gamma\gamma}<1.0$, $\Delta R_{\ell\ell}<1.4$ and $\Delta R_{\ell\gamma}>0.4$ at Cut-4.
@@
Figure 5: The distributions of the minimum distance between the leading and sub-leading photon (left) and charged lepton (right) for signal and relevant background processes at the muon collider.
In order to separate the signal from the backgrounds in the invariant mass distribution of the photon pair and the lepton pair in Fig. 6, we set Cut-5 as $110$ GeV$<m_{\gamma\gamma}<140$ GeV and $|m_{\ell\ell}-m_{Z}|<20$ GeV, where $m_{Z}=91.2$ GeV.
@@
Consequently, as illustrated in Figs. 3-6, the kinematic cuts defining deviation trends between the signal and the background and their descriptions are summarized in Table 1.
@@
Cut-0 $N_{\gamma}\geq 2$, $N_{\ell}\geq 2$ (least two same flavor opposite-sign leptons)
Cut-1 $p_{T}^{\ell_{1}}>50$ GeV, $p_{T}^{\ell_{2}}>10$ GeV, $|\eta^{\ell_{1,2}}|<2.5$
Cut-2 $p_{T}^{\gamma_{1}}>50$ GeV, $p_{T}^{\gamma_{2}}>10$ GeV, $|\eta^{\gamma_{1,2}}|<2.5$
@@
Cut-4 $\Delta R_{\gamma\gamma}<1.0$, $\Delta R_{\ell\ell}<1.4$, $\Delta R_{\ell\gamma}>0.4$
Cut-5 $110$ GeV$<m_{\gamma\gamma}<140$ GeV, $|m_{\ell\ell}-m_{Z}|<20$ GeV
The cumulative event numbers of signal ($\overline{c}_{\gamma}=1$, $\widetilde{c}_{\gamma}=1$, $\overline{c}_{HB}=1$, $\widetilde{c}_{HB}=1$, $\overline{c}_{HW}=1$, $\widetilde{c}_{HW}=1$) and relevant backgrounds ($Z\gamma\gamma$, $\tau\tau\gamma\gamma$) after each cut used in the analysis are given in Table 2. In this table, the cumulative event numbers are calculated by multiplying the cross-sections with the integrated luminosity, where the integrated luminosity of the muon collider is ${\cal L}_{\text{int}}=10$ ab-1. The signal to total background ratio ($S/B$) after each cut is presented in Table 2 to examine the efficiency of kinematic cuts in suppressing backgrounds. The impact of Cut-5 plays a significant role in suppressing relevant backgrounds by increasing the signal to total background ratio in cut flow by approximately 22 times.

### v2

@@ Abstract
We investigate the sensitivities on the Wilson coefficients of dimension-six operators associated with the anomalous $H\gamma\gamma$, $H\gamma Z$ and $HZZ$ vertices through the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ at future muon collider. Signal events involving Higgs-gauge boson interactions and relevant backgrounds events at the muon collider designed with a center-of-mass energy of 10 TeV and an integrated luminosity of 10 ab-1 are generated in Madgraph within the framework of the model-independent Standard Model effective field theory. Realistic detector effects of the muon collider detector are simulated by Delphes. The limits at 95% C.L., $3\sigma$ and $5\sigma$ on the coefficients $\overline{c}_{HB}$ and $\overline{c}_{HW}$ are obtained without and with systematic uncertainty of 15% and compared with the present experimental and phenomenological limits.
@@ I Introduction
With the discovery of the Higgs particle Aad:2012les ; Chatrchyan:2012les , intensive studies Aad:2022owv ; Tumasyan:2022uez have been done to reveal the properties of the Higgs particle at colliders, both through the Standard Model (SM) predictions and scenarios beyond the SM. Despite the discovery of the Higgs boson, the Higgs sector’s structure is still unclear. Studies Aad:2016saf ; Aad:2020czq ; Aad:2020jkn ; Sirunyan:2020wxc ; Tumasyan:2022ssx have shown that the discovered Higgs boson is a CP-even scalar with CP-conserving interactions and is consistent with SM predictions. However, the SM is not perfect and there is great motivation to search for new physics beyond the SM to address many unsolved questions.
In particle physics, one crucial area of study for future experiments is the search for CP violation. Future high-energy physics experiments have special qualifications to test for CP violation in the Higgs boson interactions. Among future collider designs, the multi-TeV muon collider incorporates many of these special features. Owing to their greater mass compared to electrons, muons emit about two billion times less synchrotron radiation, which facilitates their acceleration to high energies within a circular collider. Furthermore, as fundamental leptons rather than composite particles like protons, muons deposit the entirety of their center-of-mass energy into the collision, enhancing the effective energy reach. In addition, lepton colliders inherently offer a much cleaner experimental environment and substantially reduced background levels relative to hadron colliders, thereby improving the precision of measurements and the sensitivity to new physics phenomena. Since CP violation impacts are negligible in the SM, they offer great testing opportunities to compare the performance of future facilities. The search for precise measurements of Higgs boson couplings at multi-TeV muon colliders has recently become more important in the study of physics beyond the SM Chiesa:2020jer ; Han:2021opj ; Buttazzo:2021uha ; Han:2021egh ; Forslund:2022hjw ; Celada:2024tgv ; Chiesa:2024asx ; Cassidy:2024uhn ; Dermisek:2024qaz ; Spor:2024ghq ; Forslund:2024yhz ; Gurkanli:2025wcv ; Chen:2022ygc ; Li:2024pmz ; Spor:2024yln . The presence of theories beyond the SM that can predict effects of CP violation will have significant outcomes for particle physics in the future.
In this study, we use the Standard Model Effective Field Theory (SMEFT) as a model-independent way to parameterize the effects of new physics at the high energies of the 10 TeV muon collider to improve the search for CP-violating phenomena. Any extra CP-violating sources identified by the new physics at high energies can be expressed in terms of CP-violating SMEFT parameters in this framework Degrande:2022mcx .
@@ II Effective operators and Higgs-gauge boson interactions
A new physics model with additional fields and interactions is needed to make progress in particle physics. In order to understand the new physics in light of experimental data, it is necessary to establish the relationship between the observables at the electroweak scale and physics beyond the SM at a relatively high scale. In order to describe new physics at high mass scales, effective field theory (EFT) presents a model-independent way. EFTs are a tool to bridge the gap between scales and reveal the indirect new physics effects from experimental data. The EFT is only valid up to a particular energy scale ($\Lambda$); in order for the theory to be practical, it must be above the energy scale ($E$) that is immediately accessible experimentally. In other words, when $\Lambda\gg E$, the EFT gives a good approximation.
One of the most effective ways to measure the effects of the new physics is to use higher-dimensional effective operators consisting of SM fields that obey the SM gauge symmetry. These are named the SMEFT operators. In this approximation, additional dimension-six operators are added to the SM Lagrangian to ensure conservation of lepton and baryon numbers, and dimension-eight and higher operators are neglected for further suppression. The SMEFT operators can be expressed in a theoretical context in the Strongly Interacting Light Higgs (SILH) basis. In the EFT framework on the SILH basis, the effects of unobserved states, which are predicted to arise at energies bigger than an effective scale identified with the $W$-boson mass $m_{W}$ or with the vacuum expectation value of the Higgs field $\nu$, are parameterized using high-dimensional operators for interactions beyond the SM.
An effective theory of a light composite Higgs boson, which is responsible for electroweak symmetry breaking and arises as the pseudo-Goldstone boson from a strongly interacting sector, reveals the SILH basis Tosciri:2021yhb . The effective Lagrangian of the $H\gamma\gamma$, $HZZ$ and $H\gamma Z$ couplings in the SILH basis Giudice:2007ops is described as the sum of the Lagrangian involving the CP-conserving and CP-violating dimension-six operators and the dimension-four SM Lagrangian as follows:
@@
where ${\cal O}_{i}$ are the dimension-six operators and $\overline{c}_{i}$} are the Wilson coefficients. The effective Lagrangian with the CP-conserving and CP-violating operators corresponding to the second and third terms in Eq. (1), respectively, are shown below Giudice:2007ops ; Alloul:2014hws
@@
where $h$, $F_{\mu\nu}$ and $Z_{\mu\nu}$ are the Higgs boson field, the field strength tensors of photon and $Z$-boson, respectively. $\widetilde{F}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ and $\widetilde{Z}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}Z_{\rho\sigma}$ are the dual field strength tensors. The relationships among the Lagrangian parameters in the mass basis in Eq. (4) and the gauge basis in Eqs. (2-3) are shown below:
@@
We have performed analyses of the effects of dimension-six operators on Higgs-gauge boson couplings in MadGraph5_\_aMC@NLO Alwall:2014cvc based on Monte Carlo simulations using FeynRules Alloul:2014tfc . In the Higgs effective Lagrangian model file of the FeynRules, the SILH Lagrangian is included along with the Wilson coefficients for Higgs boson interactions.
We investigate the potential of anomalous Higgs-gauge boson couplings on the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$ at the muon collider and the sensitivity study of coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$ and $\widetilde{c}_{HW}$ in anomalous $H\gamma\gamma$, $HZZ$ and $H\gamma Z$ vertices. The stage of the muon collider with a center-of-mass energy of 10 TeV and an integrated luminosity of 10 ab-1 is considered Accettura:2023uhz . By combining high-energy collisions in a clean experimental environment, the multi-TeV muon collider creates unique potential for Higgs studies by providing both precise measurements and high-energy reach in a single machine. Therefore, a 10 TeV lepton collider that goes beyond the precision-energy dilemma is the ideal machine. Since electron-based colliders using current technology are unable to reach this energy scale, we are turning to muon colliders to probe this scale with leptons Black:2024edf .
The Feynman diagrams of the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ are shown in Fig. 1. The three diagrams in the top row are pure SM background processes, while the other six diagrams present signal processes, including new physics contributions from the anomalous $H\gamma\gamma$ (black dot), $H\gamma Z$ (red dot) and $HZZ$ (blue dot) vertices. The $H\gamma\gamma$ vertex (green circle) does not exist at the tree-level in the SM, but is present at the loop-level and contributes to the term $a_{H}$ in Eq. 8.
@@
The total cross-sections of the process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma$ as a function of coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$ and $\widetilde{c}_{HW}$ at the muon collider are given in Fig. 2. The transverse momentum $p_{T}^{\gamma}>10$ GeV and the pseudo-rapidity $|\eta^{\gamma}|<2.5$ for the photons on the final state particles are applied to determine the total cross-sections in Fig. 2. The total cross-sections in Fig. 2 were created by calculating a function of the coefficient in question with Madgraph when all other coefficients were zero. The SM cross-section including the contributions from the top three Feynman diagrams in Fig. 1 corresponds to the point $\overline{c}_{\gamma}=\overline{c}_{HB}=\overline{c}_{HW}=\widetilde{c}_{\gamma}=\widetilde{c}_{HB}=\widetilde{c}_{HW}=0$ in Fig. 2.
@@ III Analysis of Signal and Background Events
We describe the cut-based analysis and the detector simulation to determine the limits on anomalous Higgs-gauge boson couplings for the $H\gamma\gamma$, $HZZ$ and $H\gamma Z$ vertices at the muon collider. The process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$ with nonzero coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{HW}$ is considered signal process that comprises the SM contribution as well as interference between effective couplings and SM contributions. The following five relevant backgrounds are considered. (i) The SM contribution from the same final state of the signal process $\mu^{-}\mu^{+}\rightarrow Z\gamma\gamma\rightarrow\ell\ell\gamma\gamma$, where the $Z$-boson decays into a pair of charged leptons with the same flavor and opposite charges ($\ell=e,\mu$), is considered the main background process. (ii) The process $\mu^{-}\mu^{+}\rightarrow\tau\tau\gamma\gamma$ is considered a background process. (iii) Owing to the similarity in shape of the electromagnetic clusters between photons and electrons, a significant fraction of electrons can be mistakenly identified as photons. The background process $\mu^{-}\mu^{+}\rightarrow ZZ\rightarrow\ell\ell\ell\ell$, in which electrons are misidentified as photons, is considered. (iv) Another background contribution comes from the $t\bar{t}\gamma\gamma$ events, in which the top quarks decay leptonically in the $\mu^{-}\mu^{+}\rightarrow t\bar{t}\gamma\gamma$ process. (v) The background process containing jets that are misidentified as photons is located in the $\ell\ell\gamma\gamma$ signal region. In process $\mu^{-}\mu^{+}\rightarrow Zjj\rightarrow\ell\ell jj$, both photon candidates are misidentified jets.
@@
In MadGraph5_\_aMC@NLO, 500k events at $\sqrt{s}=10$ TeV are generated for different values of coefficients $\overline{c}_{\gamma}$, $\widetilde{c}_{\gamma}$, $\overline{c}_{HB}$, $\widetilde{c}_{HB}$, $\overline{c}_{HW}$, $\widetilde{c}_{HW}$ of the signal process and for the background processes in all analyses of this study. These events, which involve initial and final parton shower, fragmentation and decay, are processed through the Pythia 8.3 Bierlich:2022uzx . The Delphes 3.5.0 package Favereau:2014wfb , which allows the simulation of many multipurpose detectors, is used to simulate the realistic detector effects of the muon collider. At electron energy $E\geq 400$ GeV, the electron efficiencies are 68% for $1.95<|\eta|<2.5$, 96% for $1.22<|\eta|\leq 1.95$, 82% for $1.1<|\eta|\leq 1.22$, 96% for $0.91<|\eta|\leq 1.1$, 98% for $0.69<|\eta|\leq 0.91$ and 98% for $|\eta|\leq 0.69$. At muon energy $E\geq 2.5$ GeV, the muon efficiency is 99.9%. At photon energy $E\geq 2$ GeV, the photon efficiencies are 94% for $|\eta|<0.7$ and 90% for $0.7\leq|\eta|\leq 2.5$. In addition, many performance parameters regarding the detector simulation of the muon collider at Delphes, such as tracking efficiency, resolution, electron/muon/photon efficiency and tagging, are discussed in Ref. Selvaggi:2020yhe . Furthermore, all events are analyzed with ROOT 6 Brun:1997gqa .
@@
A first preselection is applied on events, requiring at least two photons in the $Z\gamma\gamma$ production and at least two leptons with the same flavour and opposite-charge in the decay of the $Z$-boson and this preselection is labeled with Cut-0. According to their transverse momentum, the leading and sub-leading charged leptons ($\ell_{1}$ and $\ell_{2}$) and photons ($\gamma_{1}$ and $\gamma_{2}$) are arranged as follows: $p_{T}^{\ell_{1}}>p_{T}^{\ell_{2}}$ and $p_{T}^{\gamma_{1}}>p_{T}^{\gamma_{2}}$, respectively. Fig. 3 shows the distributions of the transverse momentum of the leading photon and the leading charged lepton for the signal and the relevant background processes. It can be seen that the signal can be distinguished from the backgrounds by $p_{T}^{\gamma_{1}}>50$ GeV and $p_{T}^{\ell_{1}}>50$ GeV. The transverse momentum and pseudo-rapidity distributions of the leading and sub-leading photons for the events are required to be $p_{T}^{\gamma_{1}}>50$ GeV, $p_{T}^{\gamma_{2}}>10$ GeV and $|\eta^{\gamma_{1,2}}|\leq 2.5$, respectively, and are labeled Cut-1. However, the distributions of transverse momentum and pseudo-rapidity of the leading and sub-leading charged leptons for the events are required to be $p_{T}^{\ell_{1}}>50$ GeV, $p_{T}^{\ell_{2}}>20$ GeV and $|\eta^{\ell_{1,2}}|\leq 2.5$, respectively, and are labeled Cut-2.
@@
Figure 3: The distributions of transverse momentum of the leading photon (a), the leading charged lepton (b), the sub-leading photon (c) and the sub-leading charged lepton (d) for signal and relevant background processes at the muon collider.
@@
A clear difference between signal and background is observed in the distribution of the azimuthal angle between the leading and the sub-leading photon, the azimuthal angle between the two charged leptons in Fig. 4. The angular distance between two particles at the detector in the cylindrical coordinate system in the $\eta–\phi$ space is given by $\Delta R=\sqrt{|\Delta\eta|^{2}+|\Delta\phi|^{2}}$, where $\Delta\eta$ and $\Delta\phi$ are the pseudo-rapidity and the azimuthal angle differences, respectively. The distributions of the angular distance between the two photons, between the two charged leptons, and between the leading charged lepton and the leading photon are shown in Fig. 5. For photons and charged leptons to be well separated in phase space leading to their identification as distinct objects in the detector, kinematic cuts are needed such as $\Delta\phi_{\gamma\gamma}<0.8$ and $\Delta\phi_{\ell\ell}<1.8$ at Cut-3 and $\Delta R_{\gamma\gamma}>0.4$, $\Delta R_{\ell\ell}<1.4$ and $\Delta R_{\ell\gamma}>0.4$ at Cut-4.
@@
Figure 5: The distributions of the angular distance $\Delta R_{\gamma\gamma}$ (a), $\Delta R_{\ell\ell}$ (b) and $\Delta R_{\ell\gamma}$ (c) for signal and relevant background processes at the muon collider.
In order to separate the signal from the backgrounds in the invariant mass distribution of the photon pair and the lepton pair in Fig. 6, we set Cut-5 as 110110 $m_{\gamma\gamma}<140 GeV$ and $|m_{\ell\ell}-m_{Z}|<20$ GeV, where $m_{Z}=91.2$ GeV. The mass of the Higgs boson is approximately 125 GeV. The decay of the Higgs into two photons is a rare but distinguishable process. The invariant mass of the two photons, the products of this decay, becomes equal to the mass of the Higgs. Fig. 6(a) shows a peak in the $m_{\gamma\gamma}$ distribution around 125 GeV. This peak indicates that the Higgs boson is produced in the signal process and is critical for the experimental detection of the existence of the Higgs boson. The reason why we choose a wide range such as 110110 $m_{\gamma\gamma}<140 GeV$ instead of applying a narrow cut is to increase the contribution of $H\gamma\gamma$ vertex without completely losing the contribution of $HZ\gamma$ vertex. The mass of the $Z$-boson is approximately 91.2 GeV. Since the $Z$-boson decays to two charged leptons during the signal process, the $m_{\ell\ell}$ distribution in Fig. 6(b) shows the most events around $m_{\ell\ell}\approx 91.2$ GeV. This is called the resonance peak. This peak indicates that a real $Z$-boson was produced in the experiment and that it decays to a lepton pair. Lepton pairs from background events can have different invariant masses, but events from the $Z$-boson are heavily concentrated at 91.2 GeV. In summary, Table 1 lists all the kinematic cuts of the analysis.
@@
Cut-0 $N_{\gamma}\geq 2$, $N_{\ell}\geq 2$ (least two same flavor opposite-charge leptons)
Cut-1 $p_{T}^{\gamma_{1}}>50$ GeV, $p_{T}^{\gamma_{2}}>10$ GeV, $|\eta^{\gamma_{1,2}}|<2.5$
Cut-2 $p_{T}^{\ell_{1}}>50$ GeV, $p_{T}^{\ell_{2}}>20$ GeV, $|\eta^{\ell_{1,2}}|<2.5$
@@
Cut-4 $\Delta R_{\gamma\gamma}>0.4$, $\Delta R_{\ell\ell}<1.4$, $\Delta R_{\ell\gamma}>0.4$
Cut-5 110110 $m_{\gamma\gamma}<140 GeV$, $|m_{\ell\ell}-m_{Z}|<20$ GeV
The cumulative event numbers of the signal (in turn, each coefficient is set equal to 0.5, while all other coefficients are set to zero) and relevant backgrounds ($Z\gamma\gamma$, $\tau\tau\gamma\gamma$, $ZZ$, $t\bar{t}\gamma\gamma$, $Zjj$) after each cut used in the analysis are given in Table 2. In this table, the cumulative event numbers are calculated by multiplying the cross-sections generated in MadGraph5 passing through Pythia and using Delphes with the integrated luminosity, where the integrated luminosity of the muon collider is ${\cal L}_{\text{int}}=10$ ab-1. The signal to total background ratio ($S/B$) after each cut is presented in Table 2 to examine the efficiency of kinematic cuts in suppressing backgrounds. The impact of Cut-5 plays a significant role in suppressing relevant backgrounds by increasing the signal to total background ratio in cut flow by approximately 20 times.